OpenStudy (agentjamesbond007):

Write as a piecewise function f(x) = (4x-|x|)/x

myininaya (myininaya):

\[x>0, f(x)=\frac{4x-x}{x}=\frac{3x}{x}=3\] \[x<0, f(x)=\frac{4x-(-x)}{x}=\frac{4x+x}{x}=\frac{5x}{x}=5\] \[x=0, f(x) DNE \]